OpenStudy (anonymous):

Complete the square in the quadratic function f given by f(x) = 2x2 - 6x + 4

OpenStudy (anonymous):

are there any answer choices??

OpenStudy (anonymous):

no :/

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

well i think its x= 4/3 but i could b wrong

OpenStudy (misty1212):

is the answer supposed to look like \[f(x)=a(x-h)^2 +k\]?

OpenStudy (anonymous):

long reply holy crap lol

OpenStudy (misty1212):

that is my guess

OpenStudy (misty1212):

\[f(x)=2(x-h)^2+k\] to find \(h\) it is \(h=-\frac{b}{2a}\)

OpenStudy (misty1212):

in this example \[-\frac{b}{2a}=-\frac{-6}{2\times 2}=\frac{3}{2}\]

OpenStudy (misty1212):

\[f(x)=2(x-\frac{3}{2})^2+k\]is the first step

OpenStudy (misty1212):

to find \(k\) compute \[f(\frac{3}{2})\]